# TaoSR-AGRL: Adaptive Guided Reinforcement Learning Framework for E-commerce Search Relevance

### v1

@@ 1. Introduction
Large Language Models (LLMs)(Yang et al., 2025; DeepSeek-AI et al., 2025; Grattafiori et al., 2024) have demonstrated remarkable reasoning capabilities in complex domains like math(Havrilla et al., 2024; Zhang et al., 2024; Lai et al., 2024) and code generation(Hui et al., 2024; Seed et al., 2025). This has inspired their application to e-commerce search relevance, where methods like Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO)(Rafailov et al., 2024) are used to frame the task as a generative reasoning problem. However, these approaches often exhibit poor generalization on novel, reasoning-intensive long-tail samples, as they primarily learn to mimic patterns within the training distribution rather than acquiring robust, generalizable reasoning ability. To overcome these limitations in generalization and exploration, advanced reinforcement learning (RL) strategies like Group Relative Policy Optimization (GRPO)(Shao et al., 2024) are a promising direction. Yet, when applied to our context, their effectiveness is severely undermined by the critical challenge of reward sparsity(Cui et al., 2025; Devidze et al., 2022; Ecoffet et al., 2019). This problem reflects both the intrinsic difficulty of the task and the suboptimal policy initialization: complex long-tail queries demand long, precise chains of reasoning under multiple constraints, making rewarding trajectories inherently rare and difficult to discover; meanwhile, the over-specialization induced by SFT or DPO yields a rigid, narrow exploratory prior, drastically reducing the already low probability of reaching valid reasoning paths. The confluence of these factors leads to exploration stagnation, ultimately crippling both training efficiency and model robustness on challenging unseen instances.
@@
We conduct our research on top of our previous work TaoSR1(Dong et al., 2025), using a proprietary relevance-focused large model, Tbstar, as the base model. It employs a Mixture-of-Experts (MoE) architecture with 42 billion total parameters and 3.5 billion activated parameters. We construct a test set containing four types of challenging queries to systematically evaluate TaoSR-AGRL. Offline experiments demonstrate that TaoSR-AGRL significantly outperforms baseline methods such as DPO, and GRPO. Furthermore, online side-by-side human evaluations show that the framework achieves consistent improvements in real-world business scenarios. Experiments and ablation studies confirm the effectiveness of Rule-aware Reward Shaping and Adaptive Guided Replay in mitigating reward sparsity on long-tail samples and enhancing the model’s reasoning capabilities. The resulting model has been deployed on Taobao, delivering significant business benefits. The main contributions of this paper are as follows:
@@ 2.1. Search Relevance
The modeling of search relevance has evolved from classical statistical models like TF-IDF (Aizawa, 2003) and BM25 (Svore and Burges, 2009), which rely on lexical matching, to the Learning to Rank (LTR) paradigm. LTR introduced models such as RankNet (Burges et al., 2005) and LambdaMART (Burges, 2010), shifting from handcrafted functions to machine-learned optimization of IR metrics. The deep learning era emphasized semantic understanding over lexical overlap. After initial efforts with shallow networks like DSSM (Huang et al., 2013), the Transformer architecture (Vaswani et al., 2023) and pre-trained models like BERT (Devlin et al., 2019) established two dominant architectural patterns: the efficient bi-encoder for scalable retrieval (Karpukhin et al., 2020) and the high-precision cross-encoder for re-ranking (Nogueira and Cho, 2020), presenting distinct approaches to balance computational efficiency with model effectiveness. More recently, Large Language Models (LLMs) have initiated a paradigm shift, proposing a complementary perspective that frames relevance not merely as a matching problem, but as a reasoning task. Frameworks like LREF (Tang et al., 2025) and TaoSR1 (Dong et al., 2025) prompt LLMs to reason about e-commerce relevance, leveraging modern alignment techniques such as SFT, DPO (Rafailov et al., 2024), and GRPO (Shao et al., 2024). Despite their power, a critical limitation remains: reward sparsity (Zhang and Zuo, 2025). This issue is particularly acute in e-commerce search, as extensive alignment (SFT, DPO) for this singular task can over-specialize the model, constraining the exploratory capacity needed for GRPO to navigate the long-tail distribution. Our work, TaoSR-AGRL, introduces a novel optimization paradigm to specifically address this gap.
@@ 2.2. RLVR for LLM Reasoning
The paradigm for enhancing complex reasoning in Large Language Models (LLMs) has transitioned from Supervised Fine-Tuning (SFT) to Reinforcement Learning (RL). SFT excels at mimicking demonstrations but suffers from exposure bias and limited generalization (Bengio et al., 2015; Chu et al., 2025). In contrast, RL fosters superior generalization and exploration through direct environmental feedback (Zhang et al., 2025c). A key development is Reinforcement Learning from Verifiable Rewards (RLVR), which underpins frontier models like OpenAI-o1 (OpenAI et al., 2024) and DeepSeek-R1 (DeepSeek-AI et al., 2025). Algorithms such as GRPO use simple, outcome-based rewards to improve reasoning (Zeng et al., 2025). However, the efficacy of such methods is substantially limited in long-horizon tasks due to reward sparsity (Nan et al., 2025; Quadros et al., 2025; Tran et al., 2025). The low probability of discovering a correct solution through random exploration in such problems renders credit assignment highly challenging and hinders learning efficiency.
To address reward sparsity, prior work has explored two main avenues. The first, process supervision, densifies rewards by crediting intermediate steps. This is achieved by training Process Reward Models (PRMs) (Nye et al., 2021; Wang et al., 2023; Li et al., 2025) or using an “LLM-as-a-Judge” (Hosseini et al., 2024; Saha et al., 2025), though the latter can face robustness and consistency issues (Gu et al., 2024). The second avenue, external guidance, simplifies the learning task via hints from a superior LLM (Nath et al., 2025; Zhang et al., 2025b) or partial reasoning trajectories (Zhang et al., 2025a). Both strategies, however, incur significant computational overhead or rely on external “oracle” models. In contrast, our work differs by introducing a lightweight, on-demand guidance mechanism. It provides targeted assistance only when needed, eliminating external dependencies and fostering autonomous, efficient learning with minimal computational cost.
@@ 3.1. Data Construction
To prioritize hard cases that provide the most informative training signal, we adopt a Difficulty-Aware Dynamic Sampling strategy, following the general principle in (Yu et al., 2025). The base model performs multiple rounds of offline prediction over the training dataset, generating diverse predicted trajectories for each instance. Samples that are consistently predicted entirely incorrectly are then subjected to manual inspection, which reveals that their annotation accuracy is considerably low, indicating a substantial level of label noise in this subset. Therefore, our sampling process is designed to remove samples predicted either perfectly correctly or consistently completely incorrectly, as these extreme cases tend to be dominated by label noise or trivial patterns. The retained subset thus consists of medium-difficulty instances, which strike a balance between informativeness and annotation reliability, and are most effective for improving the model’s reasoning capacity.
@@ 3.2. Rule-aware Reward Shaping
The fidelity of the intermediate CoT is paramount for robust reasoning. However, a critical challenge arises because the aggregation mechanism of the derivation rules can mask underlying errors in the reasoning path. For example, consider a case where the correct reasoning path is “Category: Related” and “Attribute: Irrelevant”. According to the rules, this combination correctly yields a final relevance of 2-Mismatch. A flawed model, however, might correctly identify the category as “Related” but misjudge the attribute as “Mismatch”. Crucially, despite the incorrect intermediate step, the derivation rule would still lead the model to the same, correct final answer. When the model is optimized solely on the reward from this final outcome, it is incentivized to find such shortcuts rather than master the correct reasoning process. This is a classic case of reward hacking(Skalse et al., 2025), which leads to two severe consequences: (1) Impaired Generalization: The model fails to learn the genuine reasoning logic, making it brittle when encountering out-of-distribution query-item pairs; and (2) Reduced Trustworthiness and Interpretability: A low-quality reasoning chain hinders error attribution and decision traceability in practical applications.
@@
To prevent the model from earning unearned rewards on flawed or invalid trajectories, we introduce a validity gating mechanism. This gate acts as a prerequisite for all fine-grained rewards, ensuring that credit is only assigned to reasoning paths that meet two fundamental criteria: structural integrity (correct format) and conclusive correctness (correct final relevance). This design establishes a natural curriculum learning(Soviany et al., 2022). The model is first compelled to master the primary task of generating well-formed and correct final answers. Only after achieving this baseline competence does it receive the detailed, step-by-step guidance necessary to refine the faithfulness of its internal reasoning process. This staged approach efficiently channels the model’s learning, preventing it from getting lost optimizing flawed logic. This principle is formalized in our final reward function $R$:
@@ 4.1. Experimental Setup
The dataset was sampled from the online search logs of Taobao. To comprehensively evaluate the model’s performance in complex semantic scenarios, we selected four Challenging Query Categories: negation, affordable alternatives, question-answering (Q&A), and knowledge-based queries. The training data was obtained through the three-stage filtering process described in Section 3.1. In the initial filtering stage, approximately 1.7 million samples were selected from the raw logs. Subsequently, using Difficulty-Aware Dynamic Sampling, we removed samples that were either all correct or all incorrect, resulting in about 100,000 samples. From this, we further applied Undersampling for Distribution Balancing to adjust the proportions of different relevance levels, ultimately yielding a training set of approximately 40,000 samples.
@@
All experiments related to TaoSR-AGRL were implemented based on the open-source reinforcement learning framework ROLL(Wang et al., 2025). We used a training batch size of 64, a maximum input length of 2048, and a maximum generation length of 4096. The learning rate was set to 1e-6, with 16 gradient accumulation steps. For each input sample, 16 candidate responses were generated using a sampling temperature of 0.99 and a top-k of 100. For difficulty filtering, an online difficulty threshold range of [0.01, 0.9] was used. To enhance training stability, the PPO clipping range was set to $1\pm 0.2$, value function clipping to 0.5, and Advantage value clipping to $\pm 2.0$. The guidance trigger threshold $\tau$ was 0.1, and the hyperparameter weights $w_{\text{cate}}$, $w_{\text{attr}}$, and $w_{\text{reasoning}}$ were set to 0.4, 0.4, and 0.2 through preliminary experiments to balance different reward components effectively.
@@ 4.4. Ablation Study
In reinforcement learning, policy entropy is a critical measure of exploration. High entropy is essential for deep reasoning tasks, as it encourages the model to explore a diverse set of inferential strategies. We compare TaoSR-AGRL against TaoSR-AGRL-Static, a variant that applies guidance indiscriminately to all samples. As illustrated in Figure 2(c), TaoSR-AGRL consistently maintains a higher policy entropy. This suggests that excessive or non-adaptive guidance can lead to policy collapse(Juliani and Ash, 2024): the model becomes overly reliant on the guidance, causing its output distribution to become over-confident and collapse into a narrow set of familiar responses. This dependency ultimately stifles the model’s exploratory capabilities and impairs its ability to generalize.
@@ 4.5. Online Evaluation
Category GSB L34 Goodrate L23 Query Goodrate
@@
To assess the real-world effectiveness of the proposed model, we conducted large-scale online side-by-side human evaluations. TaoSR1 was used as the baseline bucket, whereas TaoSR‑AGRL was employed as the test bucket. For 2,000 authentic user queries, the top‑10 retrieval results of both systems were compared item‑by‑item. The results indicated that TaoSR‑AGRL consistently outperformed the baseline bucket, achieving average improvements of 7.11% in GSB, 1.74 pt in L34 Goodrate and 3.74 pt in L23 Query Goodrate. Gains were particularly pronounced on reasoning‑intensive tasks (e.g., Q&A and knowledge‑based queries), indicating that the relevance model effectively integrates domain‑specific e‑commerce knowledge with rule‑driven reasoning to handle complex cases.

### v2

@@ 1. Introduction
Large Language Models (LLMs)(yang2025qwen3technicalreport; deepseekai2025deepseekr1incentivizingreasoningcapability; grattafiori2024llama3herdmodels) have demonstrated remarkable reasoning capabilities in complex domains like math(havrilla2024teachinglargelanguagemodels; zhang2024llamaberrypairwiseoptimizationo1like; lai2024stepdpostepwisepreferenceoptimization) and code generation(hui2024qwen25codertechnicalreport; seed2025seedcoderletcodemodel). This has inspired their application to e-commerce search relevance, where methods like Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO)(rafailov2024directpreferenceoptimizationlanguage) are used to frame the task as a generative reasoning problem. However, these approaches often exhibit poor generalization on novel, reasoning-intensive long-tail samples, as they primarily learn to mimic patterns within the training distribution rather than acquiring robust, generalizable reasoning ability. To overcome these limitations in generalization and exploration, advanced reinforcement learning (RL) strategies like Group Relative Policy Optimization (GRPO)(shao2024deepseekmathpushinglimitsmathematical) are a promising direction. Yet, when applied to our context, their effectiveness is severely undermined by the critical challenge of reward sparsity(cui2025processreinforcementimplicitrewards; 10.5555/3600270.3600692; ecoffet2019go). This problem reflects both the intrinsic difficulty of the task and the suboptimal policy initialization: complex long-tail queries demand long, precise chains of reasoning under multiple constraints, making rewarding trajectories inherently rare and difficult to discover; meanwhile, the over-specialization induced by SFT or DPO yields a rigid, narrow exploratory prior, drastically reducing the already low probability of reaching valid reasoning paths. The confluence of these factors leads to exploration stagnation, ultimately crippling both training efficiency and model robustness on challenging unseen instances.
@@
We conduct our research on top of our previous work TaoSR1(dong2025taosr1thinkingmodelecommerce), using a proprietary relevance-focused large model, Tbstar, as the base model. It employs a Mixture-of-Experts (MoE) architecture with 42 billion total parameters and 3.5 billion activated parameters. We construct a test set containing four types of challenging queries to systematically evaluate TaoSR-AGRL. Offline experiments demonstrate that TaoSR-AGRL significantly outperforms baseline methods such as DPO, and GRPO. Furthermore, online side-by-side human evaluations show that the framework achieves consistent improvements in real-world business scenarios. Experiments and ablation studies confirm the effectiveness of Rule-aware Reward Shaping and Adaptive Guided Replay in mitigating reward sparsity on long-tail samples and enhancing the model’s reasoning capabilities. The resulting model has been deployed on Taobao, delivering significant business benefits. The main contributions of this paper are as follows:
@@ 2.1. Search Relevance
The modeling of search relevance has evolved from classical statistical models like TF-IDF (AIZAWA200345) and BM25 (10.1145/1645953.1646237), which rely on lexical matching, to the Learning to Rank (LTR) paradigm. LTR introduced models such as RankNet (10.1145/1102351.1102363) and LambdaMART (Burges2010FromRT), shifting from handcrafted functions to machine-learned optimization of IR metrics. The deep learning era emphasized semantic understanding over lexical overlap. After initial efforts with shallow networks like DSSM (huang2013learning), the Transformer architecture (vaswani2023attentionneed) and pre-trained models like BERT (devlin-etal-2019-bert) established two dominant architectural patterns: the efficient bi-encoder for scalable retrieval (karpukhin-etal-2020-dense) and the high-precision cross-encoder for re-ranking (nogueira2020passagererankingbert), presenting distinct approaches to balance computational efficiency with model effectiveness. More recently, Large Language Models (LLMs) have initiated a paradigm shift, proposing a complementary perspective that frames relevance not merely as a matching problem, but as a reasoning task. Frameworks like LREF (Tang_2025) and TaoSR1 (dong2025taosr1thinkingmodelecommerce) prompt LLMs to reason about e-commerce relevance, leveraging modern alignment techniques such as SFT, DPO (rafailov2024directpreferenceoptimizationlanguage), and GRPO (shao2024deepseekmathpushinglimitsmathematical). Despite their power, a critical limitation remains: reward sparsity (Zhang2025GRPOLEADAD). This issue is particularly acute in e-commerce search, as extensive alignment (SFT, DPO) for this singular task can over-specialize the model, constraining the exploratory capacity needed for GRPO to navigate the long-tail distribution. Our work, TaoSR-AGRL, introduces a novel optimization paradigm to specifically address this gap.
@@ 2.2. RLVR for LLM Reasoning
The paradigm for enhancing complex reasoning in Large Language Models (LLMs) has transitioned from Supervised Fine-Tuning (SFT) to Reinforcement Learning (RL). SFT excels at mimicking demonstrations but suffers from exposure bias and limited generalization (NIPS2015_e995f98d; chu2025sftmemorizesrlgeneralizes). In contrast, RL fosters superior generalization and exploration through direct environmental feedback (zhang2025onpolicyrlmeetsoffpolicy). A key development is Reinforcement Learning from Verifiable Rewards (RLVR), which underpins frontier models like OpenAI-o1 (openai2024openaio1card) and DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability). Algorithms such as GRPO use simple, outcome-based rewards to improve reasoning (zeng2025simplerlzooinvestigatingtamingzero). However, the efficacy of such methods is substantially limited in long-horizon tasks due to reward sparsity (nan2025ngrponegativeenhancedgrouprelative; quadros2025llmdrivenintrinsicmotivationsparse; tran2025exploitingtreestructurecredit). The low probability of discovering a correct solution through random exploration in such problems renders credit assignment highly challenging and hinders learning efficiency.
To address reward sparsity, prior work has explored two main avenues. The first, process supervision, densifies rewards by crediting intermediate steps. This is achieved by training Process Reward Models (PRMs) (nye2021workscratchpadsintermediatecomputation; wang2023selfconsistencyimproveschainthought; li2025enhancingreasoningprocesssupervision) or using an “LLM-as-a-Judge” (Hosseini2024VSTaRTV; Saha2025LearningTP), though the latter can face robustness and consistency issues (Gu2024ASO). The second avenue, external guidance, simplifies the learning task via hints from a superior LLM (nath2025adaptiveguidanceacceleratesreinforcement; zhang2025critiquegrpoadvancingllmreasoning) or partial reasoning trajectories (zhang2025stephintmultilevelstepwisehints). Both strategies, however, incur significant computational overhead or rely on external “oracle” models. In contrast, our work differs by introducing a lightweight, on-demand guidance mechanism. It provides targeted assistance only when needed, eliminating external dependencies and fostering autonomous, efficient learning with minimal computational cost.
@@ 3.1. Data Construction
To prioritize hard cases that provide the most informative training signal, we adopt a Difficulty-Aware Dynamic Sampling strategy, following the general principle in (yu2025dapoopensourcellmreinforcement). The base model performs multiple rounds of offline prediction over the training dataset, generating diverse predicted trajectories for each instance. Samples that are consistently predicted entirely incorrectly are then subjected to manual inspection, which reveals that their annotation accuracy is considerably low, indicating a substantial level of label noise in this subset. Therefore, our sampling process is designed to remove samples predicted either perfectly correctly or consistently completely incorrectly, as these extreme cases tend to be dominated by label noise or trivial patterns. The retained subset thus consists of medium-difficulty instances, which strike a balance between informativeness and annotation reliability, and are most effective for improving the model’s reasoning capacity.
@@ 3.2. Rule-aware Reward Shaping
The fidelity of the intermediate CoT is paramount for robust reasoning. However, a critical challenge arises because the aggregation mechanism of the derivation rules can mask underlying errors in the reasoning path. For example, consider a case where the correct reasoning path is “Category: Related” and “Attribute: Irrelevant”. According to the rules, this combination correctly yields a final relevance of 2-Mismatch. A flawed model, however, might correctly identify the category as “Related” but misjudge the attribute as “Mismatch”. Crucially, despite the incorrect intermediate step, the derivation rule would still lead the model to the same, correct final answer. When the model is optimized solely on the reward from this final outcome, it is incentivized to find such shortcuts rather than master the correct reasoning process. This is a classic case of reward hacking(skalse2025definingcharacterizingrewardhacking), which leads to two severe consequences: (1) Impaired Generalization: The model fails to learn the genuine reasoning logic, making it brittle when encountering out-of-distribution query-item pairs; and (2) Reduced Trustworthiness and Interpretability: A low-quality reasoning chain hinders error attribution and decision traceability in practical applications.
@@
To prevent the model from earning unearned rewards on flawed or invalid trajectories, we introduce a validity gating mechanism. This gate acts as a prerequisite for all fine-grained rewards, ensuring that credit is only assigned to reasoning paths that meet two fundamental criteria: structural integrity (correct format) and conclusive correctness (correct final relevance). This design establishes a natural curriculum learning(soviany2022curriculum). The model is first compelled to master the primary task of generating well-formed and correct final answers. Only after achieving this baseline competence does it receive the detailed, step-by-step guidance necessary to refine the faithfulness of its internal reasoning process. This staged approach efficiently channels the model’s learning, preventing it from getting lost optimizing flawed logic. This principle is formalized in our final reward function $R$:
@@ 4.1. Experimental Setup
The dataset was sampled from the online search logs of Taobao. To comprehensively evaluate the model’s performance in complex semantic scenarios, we selected four Challenging Query Categories: negation, affordable alternatives, question-answering (Q&A), and knowledge-based queries. The training data was obtained through the three-stage filtering process described in Section 3.1. In the initial filtering stage, millions of samples were selected from the raw logs. Subsequently, using Difficulty-Aware Dynamic Sampling, we removed samples that were either all correct or all incorrect, resulting in hundreds of thousands of samples. From this, we further applied Undersampling for Distribution Balancing to adjust the proportions of different relevance levels, ultimately yielding a training set of tens of thousands of samples.
@@
All experiments related to TaoSR-AGRL were implemented based on the open-source reinforcement learning framework ROLL(wang2025reinforcementlearningoptimizationlargescale). We used a training batch size of 64, a maximum input length of 2048, and a maximum generation length of 4096. The learning rate was set to 1e-6, with 16 gradient accumulation steps. For each input sample, 16 candidate responses were generated using a sampling temperature of 0.99 and a top-k of 100. For difficulty filtering, an online difficulty threshold range of [0.01, 0.9] was used. To enhance training stability, the PPO clipping range was set to $1\pm 0.2$, value function clipping to 0.5, and Advantage value clipping to $\pm 2.0$. The guidance trigger threshold $\tau$ was 0.1, and the hyperparameter weights $w_{\text{cate}}$, $w_{\text{attr}}$, and $w_{\text{reasoning}}$ were set to 0.4, 0.4, and 0.2 through preliminary experiments to balance different reward components effectively.
@@ 4.4. Ablation Study
In reinforcement learning, policy entropy is a critical measure of exploration. High entropy is essential for deep reasoning tasks, as it encourages the model to explore a diverse set of inferential strategies. We compare TaoSR-AGRL against TaoSR-AGRL-Static, a variant that applies guidance indiscriminately to all samples. As illustrated in Figure 2(c), TaoSR-AGRL consistently maintains a higher policy entropy. This suggests that excessive or non-adaptive guidance can lead to policy collapse(juliani2024studyplasticitylossonpolicy): the model becomes overly reliant on the guidance, causing its output distribution to become over-confident and collapse into a narrow set of familiar responses. This dependency ultimately stifles the model’s exploratory capabilities and impairs its ability to generalize.
@@ 4.5. Online Evaluation
Category GSB Query Goodrate Item Goodrate
@@
To assess the real-world effectiveness of the proposed model, we conducted large-scale online side-by-side human evaluations. TaoSR1 was used as the baseline bucket, whereas TaoSR‑AGRL was employed as the test bucket. For 2,000 authentic user queries, the top‑10 retrieval results of both systems were compared item‑by‑item. The results indicated that TaoSR‑AGRL consistently outperformed the baseline bucket, achieving average improvements of 7.11% in GSB, 1.74 pt in Query Goodrate and 3.74 pt in Item Goodrate. Gains were particularly pronounced on reasoning‑intensive tasks (e.g., Q&A and knowledge‑based queries), indicating that the relevance model effectively integrates domain‑specific e‑commerce knowledge with rule‑driven reasoning to handle complex cases.